You are jogging along at a slow speed on a cool day and are losing heat at a rate of 120 W via convection. If your surface area is 0.45 m2, your skin temperature is 26°C. and the air temperature is 17°C then what is the heat transfer coefficient of convection, hconv? (to 2 s.f and in W m−2 K−1)
1
Expert's answer
2020-05-24T18:00:36-0400

By definition, the heat transfer coefficient of convection is given by the expression (see https://en.wikipedia.org/wiki/Heat_transfer_coefficient):

"h_{conv} = \dfrac{q}{\Delta T},"

where "q = P\/A" is the heat flux (thermal power per unit area) and "\Delta T" is the difference in temperature (in Kelvin) between the body surface and air. Putting it all together:


"h_{conv} = \dfrac{P}{A\Delta T} =\dfrac{120W}{0.45m^2((273+26)K-(273+17)K)} \approx29.6 \dfrac{W}{m^2K}"

Answer. hconv = 29.6 W m-2 K-1.